a man can walk a distance 900 meters.How long will it take him to travel at rate of 8km/hr
Answer by fractalier(6550)

distance = rate times time, or d = rt.
We are looking for time t, so (for 900 m = 0.9 km)
t = d/r = .9/8 = .1125 km/hr